OpenStudy (anonymous):

if a field has perimeter p, what shape maximizes the area? How can I justify my answer?

OpenStudy (anonymous):

you could compare the ratio of Area to Perimeter for shapes. It should be fairly easy to show that the shape must be convex. Example: for an equilateral triangle \[\frac{ A }{ P }=\frac{ \frac{ 1 }{ 2 } b\cdot \frac{ b \sqrt{3} }{ 2 } }{3b }=\frac{ b \sqrt{3} }{ 12 }=\frac{b}{4\sqrt{3}}\] for a square: \[\frac{ A }{P }=\frac{ b^2 }{4b } =\frac{ b }{ 4 }\]

OpenStudy (anonymous):

it ends up being a circle but to show that you really need calculus. is this a calculus class?

OpenStudy (anonymous):

This is a problem solving math class. This is a really hard problem to figure out.

OpenStudy (anonymous):

have you taken calculus?

OpenStudy (anonymous):

I took calculus in High school and that was 5 years ago. I decided to go back to school and I have to take this class due to my major.

OpenStudy (anonymous):

so a regular polygon will maximize the area for a given perimeter, for a given polygonal shape. for example, a quadrilateral with the largest area, given a fixed perimeter, is a square. Area of a regular polygon = \(\frac{1}{2}\cdot \text{apothem} \cdot \text{perimeter}\)

OpenStudy (anonymous):

to maximize the area you must maximize the apothem. the apothem is the radius of the inscribed circle. As the number of sides increases, the shape becomes a circle and the apothem becomes the same as the radius of the circumscibed circle.

OpenStudy (anonymous):



OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

it helps a little. Thanks!

OpenStudy (anonymous):

hopefully professor can go over it wednesday.

OpenStudy (anonymous):

what part(s) are troubling you?

OpenStudy (anonymous):

first of all the word apothem